Section II, Short Answer Write your final answer in the space provided to the right. 1. Using a calculator, you find the probability of an event to be 7.7843243E-4. Express the probability in 1 decimal form rounded to the first nonzero digit. 2. For a data set, Q 42 and a Q, 70, find the value of the upper fence. 2411.5 3. Samantha took 2 classes last semester. In her 4-unit class she earned an A and in her 3-unit class she 3.GPA 91 earned a C. Find her grade point average for last semester. Round to 2 decimal places. grade points: A 4, B-3, C-2, D-1, F-o) 4. If a data set has a variance of 46, what is the standard deviation? Round to 1 decimal place Page 416
Answer # 1 & 4
Answer #1

1) probability =0.0008

2) upper fence =Q3+1.5*(Q3-Q1) =70+1.5*(70-42)=112

3)

GPA =(4*4+3*2)/(4+3)=3.14

4)

std deviation =(variance)1/2 =6.8